OpenStudy (anonymous):

the sum of four consecutive integers is -22

hero (hero):

x + x+1 + x+2 + x+3 = -22

hero (hero):

Solve for x

hero (hero):

4x + 6 = -22 4x = -28 x = -7 x + 1 = -6 x + 2 = -5 x + 3 = -4